# Optimizing Temperature for Language Models with Multi-Sample Inference

### v1

@@ B.1.7 Entropy Calculation
We define the token-level entropy $\mathcal{H}$ over a sequence of decoding steps:
$\mathcal{H}=\frac{1}{K}\sum_{t=1}^{K}\left(-\sum_{\text{j}}^{N_{0}}p_{t,\text{proper},j}\log p_{t,\text{proper},j}-\sum_{\text{j}}^{N_{1}}p_{t,\text{improper},j}\log p_{t,\text{improper},j}\right)$
@@ B.2.1 Model Hyperparameter
The toy model has the following hyperparameters:
@@
    $l_{0,i}\sim\mathcal{N}(L_{0},\sigma_{0}),\ \ l_{1,i}=L_{1}.$
  • The number of total steps $K$.
@@
The input is temperature $(T)$ and the output is the average token-level entropy $(E)$ over 500 samples. In our experiment, we set the hyperparameters to be:
@@
We test different noise tolerance rates $\alpha\in[1.5,2.0,2.5,3.0]$ to show behaviors under different noise tolerance rates.
@@ B.2.2 Result
Here is the temperature-entropy curve derived from the toy model under different noise tolerance rates $\alpha$ (Figure 8):
@@
The curves under different noise tolerances have a similar shape. Generally, the curve can be divided into two parts: (sub)-linear increase and (super)-exponential increase. In the first part, the increase is due to the various choices among the proper tokens, while the sharp rise in the second part is due to the loss of control (i.e., the model frequently chooses nonsense tokens).
The curve is very similar to the behavior of real language models, and some reference entropy curves and log-entropy curves are shown in Figure 10 and 11.
@@
It is natural to consider proper tokens can lead to correct answers and the improper tokens will result in incorrect answers, so we measure the percentage of improper tokens. As shown in Figure 9, when the temperature exceeds the turning point, the percentage of improper tokens increases quickly, implying a quality drop in samples. Interestingly, the difference in noise tolerance rates has little inference on the turning point but controls the improper token increasing speed after the turning point. Nevertheless, the percentage of improper tokens increases quickly under all tested $\alpha$.
@@ B.3 Conclusion
Our toy model provides a simplified yet insightful framework for understanding how temperature-dependent sampling dynamics may give rise to characteristic shifts in the model’s output distribution. The model predominantly chooses from the proper tokens in the low-temperature (or sub-linear growth) regime, resulting in relatively stable and controlled outputs. The distribution flattens and nonsense tokens gain significant probability mass as temperature increases beyond a certain threshold. This transition leads to a sudden and steep increase in entropy—mirroring observations in actual large language models—and a corresponding drop in the correct rate. Therefore, increasing the temperature can initially increase generation diversity (sampling among proper tokens) with a small correctness drop. However, its performance suffers a quick drop after reaching a certain threshold (i.e., the turning point).

### v2

@@ B.1.7 Entropy Calculation
We define token-level entropy $\mathcal{H}$ on a sequence of decoding steps:
$\mathcal{H}=\frac{1}{K}\sum_{t=1}^{K}\sum_{j}p_{t,j}\log p_{t,j}=-\frac{1}{K}\sum_{t=1}^{K}\left(\sum_{j=1}^{N_{0}}p_{t,\text{proper},j}\log p_{t,\text{proper},j}+\sum_{j=1}^{N_{1}}p_{t,\text{improper},j}\log p_{t,\text{improper},j}\right).$
@@ B.2.1 Model Hyperparameter
The stochastic process model has the following hyperparameters:
@@
    $l_{0,i}\sim\mathcal{N}(L_{0},\sigma_{0}),\ \ l_{1,i}=L_{1};$
  • The number of total steps $K$;
@@
The input is the temperature $T$ and the output is the average token-level entropy $\mathcal{H}$ over 500 seeds. In our experiment, the hyperparameters are set to be:
@@
Furthermore, we tested different noise tolerance rates $\alpha\in[1.5,2.0,2.5,3.0]$ to show the behavior under different noise tolerances.
@@ B.2.2 Result
Figure 8 is the temperature entropy curve derived from the toy model at different noise tolerance rates $\alpha$.
@@
The curves with different noise tolerances have a similar shape. Generally, the curve can be divided into two parts: (sub)-linear increase and (super)-exponential increase. In the first part, the increase is due to the various choices among the proper tokens, while the sharp rise in the second part is due to the loss of control (i.e., the model frequently chooses improper tokens and then makes the error rate extremely high).
In particular, the curve is very similar to the behavior of real language models, and some reference entropy curves and log-entropy curves are shown in Figures 10 and 11.
@@
It is natural to consider that proper tokens can lead to correct final answers and that improper tokens will result in incorrect final answers, so we measure the percentage of improper tokens. As shown in Figure 9, when the temperature exceeds the turning point, the percentage of improper tokens increases quickly, implying a quality drop in the samples. Interestingly, the difference in noise tolerance rates has little inference on the turning point but controls the improper token increasing speed after the turning point. However, the percentage of improper tokens increases rapidly in all tested $\alpha$.
@@ B.3 Conclusion
Our stochastic process model provides a simplified but insightful framework for understanding how temperature-dependent sampling dynamics can lead to characteristic shifts in the model output distribution. The model predominantly chooses from the proper tokens in the low-temperature (or sublinear growth) regime, resulting in relatively stable and controlled outputs. The distribution flattens, and nonsense tokens gain significant probability mass as temperature increases beyond a certain threshold. This transition leads to a sudden and steep increase in entropy—mirroring observations in actual language models—and a corresponding drop in the correct rate. Therefore, increasing the temperature can initially increase generation diversity (sampling among proper tokens) with a small correctness drop. However, the performance suffers a quick drop after reaching a certain threshold (i.e., the turning point EntP).